OpenStudy (anonymous):

Nisha closes the nozzle of the funnel and fills it completely with a liquid. She then opens the nozzle. If the liquid drips at the rate of 15 cubic inches per minute, how long will it take for all the liquid to pass through the nozzle? (Use π = 3.14.) (1 points) 3.14 minutes 5.02 minutes 2.51 minutes 3.76 minutes

OpenStudy (elonasushchik):

its the second one @grex10

OpenStudy (elonasushchik):

First you need to find the volume of the cone

OpenStudy (elonasushchik):

The volume of any cone with radius 'r' and height h is V = (pi*r^2*h)/3

OpenStudy (elonasushchik):

which is 75.36

OpenStudy (elonasushchik):

Once you know the volume of the cone, you can use it to figure out how long that amount of liquid will take to drip out of the cone

OpenStudy (elonasushchik):

Now we have this idea Amount dripped out = (Rate of drip)*(Time)

OpenStudy (elonasushchik):

So this means Amount dripped out = (Rate of drip)*(Time) 75.36 = 15*t where t is the time Note: we set "Amount dripped out" to all 75.36 cubic inches because want all the liquid to drip out

OpenStudy (elonasushchik):

So your last step is to solve for t which is 5.024

OpenStudy (elonasushchik):

@grex10 U there....I explained it tell me if u like understand what I wrote

OpenStudy (anonymous):

I'm a bit confused